OpenStudy (anonymous):

Help me integrate \[\int\limits_{}^{} \sin^5 3x \cos ^3 3x dx\]

OpenStudy (watchmath):

Use \(u=\sin (3x)\). Then \(du=3\cos(3x)\,dx\) The integral become \(\frac{1}{3}\int u^5\cos^2(3x)\,du=(1/3)\int u^5(1-u^2)\, du\)

OpenStudy (anonymous):

Wow thank you!